OpenStudy (weasel123):

is 2 √x+5+7=4 extraneous? will give medal

zepdrix (zepdrix):

\[\large\rm 2\sqrt{x+5}+7=4\]For these types of problems, try to solve for the square root.

zepdrix (zepdrix):

Subtracting 7 from each side,\[\large\rm 2\sqrt{x+5}=-3\]Dividing by 2,\[\large\rm \sqrt{x+5}=-\frac32\]

zepdrix (zepdrix):

Square root operation should always produce a `positive number`. There is no value that when we take the root of it, gives us a -3/2.

OpenStudy (weasel123):

so is it?